OpenStudy (anonymous):

How do you use the half angle identity for cos (-pi/8) ?

pooja195 (pooja195):

Cos(Pi/8)=[(1+Cos{Pi/4})/2]^1/2 Cos(Pi/4)=SQRT2/2=.707 Therefore Cos(Pi/8)=[(1+.707)/2]^1/2=.924